a spherical balloon is inflated at a rate of 36pi cm³/sec. How fast is the radius of the balloon increasing when radius is 7cm?

1
Expert's answer
2021-12-01T17:15:17-0500

"V=\frac{4}{3}\pi r^3."

"\frac{dV}{dt}=4\pi r^2\frac{dr}{dt}."

"\frac{dr}{dt}=\frac{\frac{dV}{dt}}{4\pi r^2}"

When "r=7:" "\frac{dr}{dt}=\frac{36}{4\pi *7^2}=\frac{9}{49\pi}\approx0.0585\;\frac{cm}{sec}."